OpenStudy (mikey):

plz help for medal Divide x to the 1 half power divided by x to the 2 sevenths power. x to the 2 fourteenths power x to the 3 fourteenths power x to the 11 fourteenths power x to the 1 fifth power

OpenStudy (mikey):

@3mar

OpenStudy (jiteshmeghwal9):

\[x^{\frac{1}{2}}*x^{\frac{2}{7}}\]

OpenStudy (jiteshmeghwal9):

U must remember the property \[a^b*a^c=a^{b+c}\]

OpenStudy (mikey):

i dont understand it

OpenStudy (jiteshmeghwal9):

\[x^{\frac{1}{2}}*x^{\frac{2}{7}}\] when bases of two exponents are common and they are multiplied then their powers get added viz. \[x^{\frac{1}{2}+\frac{2}{7}}\]

OpenStudy (mikey):

x^1/3

OpenStudy (jiteshmeghwal9):

No

OpenStudy (jiteshmeghwal9):

Try again all that is remaining is adding of fractions in power

OpenStudy (johnweldon1993):

For multiplication exponent rule to apply, we would need to write this as \(\large x^{1/2}*x^{-2/7}\) Then we can continue For division we have a similar rule: \(\large \frac{x^a}{x^b} = x^{a-b}\) So here we have \(\large \frac{x^{1/2}}{x^{2/7}}=?\)

OpenStudy (mathmale):

Given "Divide x to the 1 half power divided by x to the 2 sevenths power." I would immediately write that symbolically and ditch the verbal expression. What is the rule of exponents pertaining to the problem at hand? You will need to subtract the exponent found in the denominator (2/7) from that found in the numerator (1/2). This subtraction will require identifying and using the LCD.